OpenStudy (anonymous):

A standard showerhead in Sonia's house dispenses 12 gallons of water per minute. Sonia changed this showerhead to an energy-saving one. The equation shows the amount of water dispensed, y, as a function of the number of minutes, x, for the new showerhead. y = 3x How much water does Sonia save each day with the change in showerheads if she uses the shower for 8 minutes each day? 9 gallons 17 gallons 72 gallons 93 gallons

OpenStudy (anonymous):

please help me

OpenStudy (danjs):

The old shower uses \[\frac{ 12 ~gal }{ 1~\min }*\frac{ 8~\min }{ 1 ~day } = \frac{ 12*8 ~gal }{ day }\]

OpenStudy (anonymous):

ok

OpenStudy (danjs):

Use the function to find the gallons per day for the new one... 8 min per day = x Gallons used by new shower = 3 * x

OpenStudy (danjs):

So the new one uses 25 gallons each day if used for 8 min per day ... the difference between the old and the new is the answer they want

OpenStudy (danjs):

24 gallons, sorry , not 25

OpenStudy (anonymous):

ok i understand it now

OpenStudy (danjs):

Water saved each day = (Old shower rate) - (New Shower Rate) =

OpenStudy (danjs):

k

OpenStudy (anonymous):

wait

OpenStudy (danjs):

ya

OpenStudy (anonymous):

24-12=12 but 12 isnt one of the answer choices

OpenStudy (danjs):

Old shower rate = (12*8) = 96 gal/day New Shower Rate = 3*8 = 24 gal/day

OpenStudy (anonymous):

oooooooooooooooooh

OpenStudy (anonymous):

so its 96-24

OpenStudy (danjs):

Yes

OpenStudy (anonymous):

ok thanks

OpenStudy (danjs):

If you dont know to multiply or divide, Look at the first thing i typed and notice how the units cancel to give you what you are looking for, gal/day

OpenStudy (danjs):

welcome